OpenStudy (anonymous):

integration of (sin^5x)*(cos^3x)dx

hartnn (hartnn):

split cos^3 x = cos x * cos^2 x write cos^2 x = 1- sin^2 x then its easy to see that u = sin x du =cos x dx.

OpenStudy (anonymous):

so do you leave sin^5x as it is

hartnn (hartnn):

yes, then you'll use it... s^5 (1-s^2) c dx [s=sin, c=cos]

hartnn (hartnn):

s^5 (1-s^2) c dx = (s^5 -s^7) cdx

OpenStudy (anonymous):

ah i see how that tricked me i tried to think it was a half angle formula or something

hartnn (hartnn):

so, now can you solve that integral ?

OpenStudy (anonymous):

yea i can thanks a lot bro

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

math not my best subject :P

hartnn (hartnn):

it became my best subject, after lot of practicing :)

OpenStudy (anonymous):

1 question though can you split sin of x to solve also

hartnn (hartnn):

yes,but that will be more complicated....but yes.

OpenStudy (anonymous):

ah i see ty again

OpenStudy (anonymous):

if i need help again asking you :P